CRY+VIS Interface Meeting
Date: 2021/9/17 13:00-13:35
Participants: Takahashi, Kimura, Hirata, Sato, Yokozawa, Ushiba, Yasui, Ikeda, Akutsu, Node
- IX: Health check. Investigation of the disturbance. BF damper itself doesn't cause. Diagonalization of IP was not good. Effect due to the convection will be checked.
- EX: Health check. Gurdian is developed. Installation of ACCs will be scheduled after the health check.
- IY: Some cables were repaired.
- EY: Adjustable height range was measured. F0 FR should be modified.
- TM height setting: Summary of adjustable height range.
- Commissioning in air: Health check.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Unused ports will be closed.
- Heater control: Implemented to BS and SRs.
Date: 2021/9/10 14:00-14:50
Participants: Takahashi, Ushiba, Hirata, Sato, Aso, Fabian, Terrence, Ishizaki, Tanaka
- Reset the position of limit switch for the F0 vertical FR in IX.
- Measured the gap between the damper magnets and the Cu plate in IX.
- Checked the adjustable height range in EY.
- This week lots of people worked for Type-B.
On Monday we double-checked the problems that were detected during the cable health check (klog 18237).
- The three important ones are:
- BS F1 LVDT primary coil has a connection to the chamber.
- SRM TM-H1 coil has a connection to the chamber.
- SR2 IM-V3 OSEM coil has a resistance higher than expected.
- Others are related to F0 yaw stepper motors and the SRM IP LVDTs, but these devices are accessible and can be fixed easily.
On Tuesday and Wednesday we checked the BS F1 LVDT: after various tests we determined the problem is at the level of the F1 Filter, which is not accessible. What are our options?
Take the suspension out and check: possible, but time consuming and risky, especially because the problem may arise inside of the F1 filter, situation in which we wouldn't be able to fix the problem without dissasembling the system.
Modify the LVDT driver from a differential reference signal output to a single-ended one: possible I guess.
- On Thursday we raised the SR2 and SR3 pre-isolators using the new height markings made by Tomaru-san and Yasui-san: the only highlight was that we had to leave SR3 mirror 0.85 mm (approx.) below the target because one of the jacks reached its higher limit.
- On Friday we checked the SRM TM-H1 coil (klog pending):
- Problem: pin 7 is connected to the chamber and to pin 5. The coil is between pins 2 and 7.
After various tests, we determined that the cable going from the pre-isolator to the BF is the culprit.
- We removed all the aluminiun foil from the connectors in the line, but pin 5 still has a connection to ground! We are puzzled by this!
- A good solution would be to use pins 1 and 6 in that cable instead of 2 and 7.
- I dropped an M4 screw on top of the F1 filter. We'll look for it.
On Monday we double-checked the problems that were detected during the cable health check (klog 18240): we need to check PR3 BF LVDTs H3 and V3, which have higher resistances than expected in one of their coils.
- Type-A team: Health check in IX and EX.
- Type-B team: Cable check/repair.
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
Effectiveness of ”BS outer frame" reinforcement. BS outer frame.pdf
The week from 9/20
- Hirata: 9/21-22
- Yano: 9/21-22
The week from 9/27
- Takahashi: 9/27-10/1
- Sato: 9/27-10/1
- Hirata: 9/27-10/1
- Yano: 9/27-10/1